OpenStudy (anonymous):

A rectangle has an area of 16 square feet. What is its maximum perimeter?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

no, not 16, if x=16 and y=1 perimeter is 16+16+1+1=34 which is more than 16, maximum perimeter is infinity i think when x is very big and y is very small

OpenStudy (anonymous):

it is maximum and it is for a square and for rectangles it sud be less than 16

OpenStudy (anonymous):

i said 3.

OpenStudy (anonymous):

for rectangle it should be Area = 16ft^2 x*y = 16ft^2 now find out x & y

OpenStudy (anonymous):

so 16 is not maximum perimeter, because 16 is only when x and y are equal to 4, but if x=16 and y =1? what's perimeter then?

OpenStudy (anonymous):

the p=34 i think

OpenStudy (anonymous):

yeah thats true

OpenStudy (anonymous):

maximum perimeter is infinity..........

OpenStudy (anonymous):

so yeah as i said before that it's infinity, and it's probably mistake in question, it should have asked for minimum perimeter

jhonyy9 (jhonyy9):

so than A=16 we can getting this from 4*4 or 2*8 or 16*1- just using integers - hence we get the greater perimeter using sides dimensions 16 and 1 because than we get 16+16+1+1=34

OpenStudy (anonymous):

it doesn't say that we can use only integers